# Does Refusal Training in LLMs Generalize to the Past Tense?

### v1

@@ Abstract
Refusal training is widely used to prevent LLMs from generating harmful, undesirable, or illegal outputs. We reveal a curious generalization gap in the current refusal training approaches: simply reformulating a harmful request in the past tense (e.g., ”How to make a Molotov cocktail?” to ”How did people make a Molotov cocktail?”) is often sufficient to jailbreak many state-of-the-art LLMs. We systematically evaluate this method on Llama-3 8B, GPT-3.5 Turbo, Gemma-2 9B, Phi-3-Mini, GPT-4o, and R2D2 models using GPT-3.5 Turbo as a reformulation model. For example, the success rate of this simple attack on GPT-4o increases from 1% using direct requests to 88% using 20 past tense reformulation attempts on harmful requests from JailbreakBench with GPT-4 as a jailbreak judge. Interestingly, we also find that reformulations in the future tense are less effective, suggesting that refusal guardrails tend to consider past historical questions more benign than hypothetical future questions. Moreover, our experiments on fine-tuning GPT-3.5 Turbo show that defending against past reformulations is feasible when past tense examples are explicitly included in the fine-tuning data. Overall, our findings highlight that the widely used alignment techniques—such as SFT, RLHF, and adversarial training—employed to align the studied models can be brittle and do not always generalize as intended. We provide code and jailbreak artifacts at https://github.com/tml-epfl/llm-past-tense.
@@ 1 Introduction
  • We show that past tense reformulations lead to a surprisingly effective attack on many recent leading LLMs. We show quantitative results on Llama-3 8B, GPT-3.5 Turbo, Gemma-2 9B, Phi-3-Mini, GPT-4o, and R2D2 in Table 1 and qualitative examples on GPT-4o in Figure 1.
@@ 2 Methodology
We evaluate six target LLMs: Llama-3 8B (AI@Meta, 2024), GPT-3.5 Turbo (OpenAI, 2023), Gemma-2 9B (DeepMind, 2024), Phi-3-Mini (Abdin et al., 2024), GPT-4o (OpenAI, 2024), and R2D2 (Mazeika et al., 2024). Most of these models use supervised fine-tuning and reinforcement learning from human feedback for refusal training. In addition, R2D2 uses adversarial training against the GCG attack on top of SFT and DPO used to fine-tune the Zephyr model (Tunstall et al., 2023). For Llama-3 8B, we use the refusal-enhancing prompt introduced in Llama-2 (Touvron et al., 2023), while for the rest of the LLMs, we use their default system prompts. We list all system prompts in Appendix A.
@@ 3 Systematic Evaluation of the Past Tense Attack
We present our main results in Table 1, which show that the past tense attack performs surprisingly well, even against the most recent LLMs such as GPT-4o and Phi-3, and in many cases is sufficient to circumvent built-in safety mechanisms. For example, the attack success rate (ASR) on GPT-4o increases from 1% using direct requests to 88% using 20 past tense reformulation attempts according to the GPT-4 judge. The Llama-3 70B and rule-based judge also indicate a high ASR, although slightly lower, at 65% and 73% respectively. Similarly, evaluation on other models indicates a high ASR: for Phi-3-Mini it increases from 6% to 82%, and for R2D2, it increases from 23% to 98%. Interestingly, GPT-3.5 Turbo is slightly more robust to past tense reformulations than GPT-4o, with a 74% ASR compared to 88% for GPT-4o. To compare these numbers with established methods, we evaluate the transfer of request-specific GCG suffixes from Chao et al. (2024) optimized on Vicuna. In the same evaluation setting, these suffixes result in a 47% ASR for GPT-3.5 Turbo and only a 1% ASR for GPT-4o, according to the Llama-3 70B judge. This discrepancy shows how later iterations of frontier LLMs can patch known attacks, but remain vulnerable to new ones. Additionally, we plot the ASR over the 20 attempts in Figure 2 for all models and judges. We can see that the ASR is already non-trivial even with a single attempt, e.g., 57% success rate on GPT-4o, which is in contrast with only 1% ASR with a direct request in the present tense. Moreover, the ASR often begins to saturate after 10 attempts, which justifies our choice of 20 attempts in total.
@@
Figure 3: Attack success rate for all models according to GPT-4 as a judge across 10 harmful categories of JBB-Behaviors.
@@
It is natural to ask whether the past tense is particularly important or if the future tense works equally well. We repeat the same experiment, this time asking GPT-3.5 Turbo to reformulate requests in the future tense, using the prompt shown in Table 9. We present the results in Table 3, which shows that future tense reformulations are less effective but still have much higher ASR than direct requests. This outcome prompts two potential hypotheses: (a) The fine-tuning datasets may contain a higher proportion of harmful requests expressed in the future tense or as hypothetical events. (b) The model’s internal reasoning might interpret future-oriented requests as potentially more harmful, whereas past-tense statements, such as historical events, could be perceived as more benign.
@@ Acknowledgements
We thank Francesco Croce for many insightful discussions and feedback on a draft of this paper. We are grateful to OpenAI for providing us API credits as part of the OpenAI Researcher Access Program. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship. This work was supported by the Swiss National Science Foundation (grant number 212111).

### v2

@@ Abstract
Refusal training is widely used to prevent LLMs from generating harmful, undesirable, or illegal outputs. We reveal a curious generalization gap in the current refusal training approaches: simply reformulating a harmful request in the past tense (e.g., ”How to make a Molotov cocktail?” to ”How did people make a Molotov cocktail?”) is often sufficient to jailbreak many state-of-the-art LLMs. We systematically evaluate this method on Llama-3 8B, Claude-3.5 Sonnet, GPT-3.5 Turbo, Gemma-2 9B, Phi-3-Mini, GPT-4o-mini, GPT-4o, and R2D2 models using GPT-3.5 Turbo as a reformulation model. For example, the success rate of this simple attack on GPT-4o increases from 1% using direct requests to 88% using 20 past tense reformulation attempts on harmful requests from JailbreakBench with GPT-4 as a jailbreak judge. Interestingly, we also find that reformulations in the future tense are less effective, suggesting that refusal guardrails tend to consider past historical questions more benign than hypothetical future questions. Moreover, our experiments on fine-tuning GPT-3.5 Turbo show that defending against past reformulations is feasible when past tense examples are explicitly included in the fine-tuning data. Overall, our findings highlight that the widely used alignment techniques—such as SFT, RLHF, and adversarial training—employed to align the studied models can be brittle and do not always generalize as intended. We provide code and jailbreak artifacts at https://github.com/tml-epfl/llm-past-tense.
@@ 1 Introduction
  • We show that past tense reformulations lead to a surprisingly effective attack on many recent leading LLMs. We show quantitative results on Llama-3 8B, Claude-3.5 Sonnet, GPT-3.5 Turbo, Gemma-2 9B, Phi-3-Mini, GPT-4o mini, GPT-4o, and R2D2 in Table 1 and qualitative examples on GPT-4o in Figure 1.
@@ 2 Methodology
We evaluate eight target LLMs: Llama-3 8B (AI@Meta, 2024), Claude-3.5 Sonnet (Anthropic, 2024), GPT-3.5 Turbo (OpenAI, 2023), Gemma-2 9B (DeepMind, 2024), Phi-3-Mini (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), and R2D2 (Mazeika et al., 2024). Most of these models use supervised fine-tuning and reinforcement learning from human feedback for refusal training. In addition, R2D2 uses adversarial training against the GCG attack on top of SFT and DPO used to fine-tune the Zephyr model (Tunstall et al., 2023). For Llama-3 8B, we use the refusal-enhancing prompt introduced in Llama-2 (Touvron et al., 2023), while for the rest of the LLMs, we use their default system prompts. We list all system prompts in Appendix A.
@@ 3 Systematic Evaluation of the Past Tense Attack
We present our main results in Table 1, which show that the past tense attack performs surprisingly well, even against the most recent LLMs such as Claude-3.5 Sonnet, GPT-4o, and Phi-3-Mini, and in many cases is sufficient to circumvent built-in safety mechanisms. For example, the attack success rate (ASR) on GPT-4o mini and GPT-4o increases from 1% using direct requests to 83% and 88% using 20 past tense reformulation attempts according to the GPT-4 judge. The similarity in the numbers between these two models suggests that they were likely aligned using the same methodology. The Llama-3 70B and rule-based judge also indicate a high ASR on GPT-4o, although slightly lower, at 65% and 73% respectively. Similarly, evaluation on other models indicates a high ASR: for Claude-3.5 Sonnet, it increases from 0% to 53%, for Phi-3-Mini it increases from 6% to 82%, and for R2D2, it increases from 23% to 98%. Interestingly, GPT-3.5 Turbo is slightly more robust to past tense reformulations than GPT-4o, with a 74% ASR compared to 88% for GPT-4o. To compare these numbers with established methods, we evaluate the transfer of request-specific GCG suffixes from Chao et al. (2024) optimized on Vicuna. In the same evaluation setting, these suffixes result in a 47% ASR for GPT-3.5 Turbo and only a 1% ASR for GPT-4o, according to the Llama-3 70B judge. This discrepancy shows how later iterations of frontier LLMs can patch known attacks, but remain vulnerable to new ones. Additionally, we plot the ASR over the 20 attempts in Figure 2 for all models and judges. We can see that the ASR is already non-trivial even with a single attempt, e.g., 57% success rate on GPT-4o, which is in contrast with only 1% ASR with a direct request in the present tense. Moreover, the ASR often begins to saturate after 10 attempts, which justifies our choice of 20 attempts in total.
@@
Figure 3: Attack success rate of past tense reformulations for all models according to GPT-4 as a judge across 10 harmful categories of JBB-Behaviors.
@@
It is natural to ask whether the past tense is particularly important or if the future tense works equally well. We repeat the same experiment, this time asking GPT-3.5 Turbo to reformulate requests in the future tense, using the prompt shown in Table 9. We present the results in Table 3, which shows that future tense reformulations are less effective but still have much higher ASR than direct requests, except Claude-3.5 Sonnet on which the ASR is only 5%. This outcome prompts two potential hypotheses: (a) The fine-tuning datasets may contain a higher proportion of harmful requests expressed in the future tense or as hypothetical events. (b) The model’s internal reasoning might interpret future-oriented requests as potentially more harmful, whereas past-tense statements, such as historical events, could be perceived as more benign.
@@ Acknowledgements
We thank Francesco Croce for many insightful discussions and feedback on a draft of this paper. We are grateful to OpenAI for providing us API credits as part of the OpenAI Researcher Access Program. We thank Sudarsh Kunnavakkam for providing the API key to run the experiments on Claude-3.5 Sonnet. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship. This work was supported by the Swiss National Science Foundation (grant number 212111).